OpenStudy (anonymous):

Find f(x) if f'(x) = 3cos(x/4) and f(pi) = 9 sqrt(2)

OpenStudy (anonymous):

did you find the anti derivative?

OpenStudy (anonymous):

i know the answer: \[3\sqrt{2} + 12 \sin (\frac{x}{8}) + c\] but i only went as far as: \[12 \sin (\frac{x}{8})\] and I can't figure out where the \[3\sqrt2\] came from

OpenStudy (kainui):

Are you sure that's the answer?

OpenStudy (kainui):

So if you want f(x) given f'(x) then you know it within a constant of integration. f(x)=12sin(x/4)+C not divided by 8 on the inside. From there you can plug in x=pi and f(pi)=9sqrt(2) to solve for C.

OpenStudy (anonymous):

tried that but it didn' work out where does the \[3\sqrt2\] come from thats the problem